# Modular degree and a conjecture of Watkins

### v2

@@ 3.3. Elliptic curves with prime conductor
Let $E/\mathbb{Q}$ be any elliptic curve of positive rank, Mestre’s result in [28] implies that $m_{E}=\langle v_{E},v_{E}\rangle=\sum_{i=1}^{n}w_{i}v_{E}(e_{i})^{2}$. On the other hand, it turns out that $\sum_{i=1}^{n}v_{E}(e_{i})=0$, and all the $w_{i}$ are $1$ when $p=1\pmod{4}$. We immediately have $m_{E}$ even in this case. For the case $p=3\pmod{4}$, This settles down the conjecture for $\mathrm{rank}(E(\mathbb{Q}))=1$ case. For rank $2$ case, Gross-Kudla’s formula implies $\sum_{i=1}^{n}v_{E}(e_{i})^{3}=0$. The reader may take a look at [25] for more details. Consequently we have $\sum_{i=1}^{n}v_{E}(e_{i})^{2}=0\pmod{4}$, if each of the odd terms in $\{V_{E}(e_{i})\}_{1\leq i\leq n}$ appears an even number of times. This is literally true, provided that the following conjecture holds.
@@
When the root number of $E$ is $-1$, Kazalicki and Kohen showed that all such $\nu(e_{i})$ are, in fact, 00. This is true because the root number condition implies $t_{p}(v_{E})=-v_{E}$. For each $1\leq i\leq n$, set $1\leq\bar{i}\leq n$ be the unique index such that $E_{\bar{i}}=E^{p}_{i}$. Note that, $i=\bar{i}$ if and only if, $j(E_{i})\in\mathbb{F}_{p}$. With this, we have [24] that $t_{p}(e_{i})=e_{\bar{i}}$, which gives a proof of the conjecture for root number $-1$ case. In the root number $1$ case, the conjecture is known to be true if $E$ has a positive discriminant and there exists a prime $\ell$ such that
(9) $\left(\frac{-p}{\ell}\right)=-1,~{}a(\ell)=1\pmod{2}.$
Under these conditions, we have $B_{i,j}(\ell)=0\pmod{2}$, whenever $i,j\in S_{p}$. Moreover, the conditions in (9) are satisfied for any $E/\mathbb{Q}$ of conductor $p$, satisfying $E(\mathbb{Q})[2]$ is trivial and $\Delta_{E}>0$.
The main ingredients are $B_{i,j}(\ell)=0\pmod{2}$ and $a(\ell)=1\pmod{2}$. In the case of $\Delta_{E}<0$, the condition $\left(\frac{-p}{\ell}\right)=-1$ implies $\ell$ is inert in $\mathbb{Q}(\sqrt{-p})$ which implies that $2$ divides the order of $\mathrm{Frob}_{\ell}$, which implies that $a(\ell)$ is not odd. This is one of the main hurdles in the case of root number $1$ and negative discriminant. However, Kazalicki and Kohen’s proof also shows that it is enough to have some integer $m$ for which $a(m)=1\pmod{2}$ and all the $B_{i,j}(m)$ are even. From the commutativity of Brandt’s matrices, and [24, Proposition 20] it follows that $\widetilde{B}(m)=\left(B_{i,j}(n)\right)_{i,j\in S_{p}}\pmod{2}$ is also commutative inside $M_{s_{p}}(\mathbb{Z}/2\mathbb{Z})$. Therefore it may be fruitful to analyze the structure of those $\widetilde{B}(m)$, whose one of the eigenvalues is $1$. Also, note that each $B_{i,j}(m)$ gives a modular form $\theta_{i,j}$, from which we get another modular form $\theta^{\prime}_{i,j}:=\theta_{i,j}-E_{2}$, where $E_{2}=1-24\sum_{n=1}^{\infty}\frac{nq^{n}}{1-q^{n}}$. Now we require the cusp forms $f_{E}$ and $\theta^{\prime}_{i,j}$ to have a common Fourier coefficient of opposite parity. A statement about the independency of the Fourier coefficients, e.g., Generalized Sato-Tate hypothesis [5, Section 4.1], could be helpful in this approach.
We end the discussion of this section with the following remark. As long as the prime conductors are concerned, this is predicted by Watkins in [40, Section 4.2] the following.
@@ 3.3.1. A conjecture on the supersingular coefficients
In this section, we discuss the Conjecture 2. In [24], Kazalicki and Kohen studied modular forms over $\overline{\mathbb{F}_{p}}$ in the sense of Katz’s modular form. If we are a bit more specific, there exists a polynomial $P_{E}(x)\in\mathbb{Z}[x]$ whose roots in $\mathbb{F}_{p}$ are often supersingular zeros. In particular, they showed that if $j\in\mathbb{F}_{p}$ is $j$ invariant of a supersingular elliptic curve $E_{i}$ over $\overline{\mathbb{F}}_{p}$, then $P_{E}(j)=0\pmod{p}$ if and only if, $v_{E}(e_{i})=0\pmod{p}$. In the sense of Katz, the condition $P_{E}(j)=0\pmod{p}$ corresponds to a rule that is 00 acting on $E_{i}$. Finally, a study of Hecke operators on the space of modular forms corresponding to supersingular elliptic curves, again in the sense of Katz. This, of course, does not settle the conjecture; however, this shows that in the case of root number $-1$ case, any supersingular $j$ invariant in $\mathbb{F}_{p}$ is a root of $P_{E}(x)\pmod{p}$. We denote $s_{p}$ to be #$\#S_{p}$, i.e., the number of isomorphism classes of supersingular elliptic curves over $\mathbb{F}_{p}$. Following a remark in Kazalicki-Kohen in [24], we can deduce the following in support of the conjecture.
Theorem 5.
@@
Before proving this, let us first define the required terminology and recall the preliminary results related to the supersingular elliptic curves. For any integer $D=3\pmod{4}$, we say that an elliptic curve over any field $F$ has CM over $\mathcal{O}_{-D}=\mathbb{Z}[\frac{1}{2}(D+\sqrt{-D})]$ if, $\mathcal{O}_{-D}$ contains maximally in the endomorphism ring of the elliptic curve over $\overline{F}$. Then supersingular elliptic curves with $j$-invariant in $\mathbb{F}_{p}$ correspond to those $D$ with those $D$ for which $p$ is ramified or inert in $\mathbb{Q}(\sqrt{-D})$. On the other hand, there are $h(-D)$ many $\overline{\mathbb{Q}}$ isomorphism classes of elliptic curves over $\overline{\mathbb{Q}}$ with CM by $\mathcal{O}_{-D}$, where $h(-D)$ is the class number of $\mathbb{Q}(\sqrt{-D})$. Furthermore, we denote $h_{i}(-D)$ to be the number of optimal embeddings of the order of discriminant $-D$ into $\mathrm{End}(E_{i})$ modulo conjugation by $\mathrm{End}(E_{i})^{*}$ and $u(-D)$ is the number of units of the order. Then we set $b_{D}=\frac{1}{u(-D)}\sum_{i=1}^{n}h_{i}(-D)v_{E}(e_{i}).$
Proof of Theorem 5.
@@
This shows that if there is one even $v_{E}(e_{i})$, then there are at least two even terms since $s_{p}$ is even. Since $\mathrm{rank}(E(\mathbb{Q}))>0$, we have $L(E,1)=0$. In particular, by Gross-Waldspurger formula [19], we have $\langle v_{E},b_{D}\rangle=0$ for any $D>0$ with $\left(\frac{-D}{p}\right)=-1$.
For proof of part (a), note that since $p$ and $D$ satisfy the imposed conditions, $p$ is inert in $\mathbb{Q}(\sqrt{-D})$. In particular, $b_{D}$ is supported only on the indices in $S_{p}$. Moreover, since $\mathbb{Q}(\sqrt{-D})$ has class number $1$, we have $h_{i}(D)=1$ for exactly one $i\in S_{p}$, and that $E_{i}$ has CM by $\mathcal{O}_{-D}$. Then the condition $\langle v_{E},b_{D}\rangle=0$ implies $v_{E}(e_{i})=0$. In particular, if $s_{p}=2$, then then $v_{E}(e_{i^{\prime}})=0$ for the other $i^{\prime}\in S_{p}$, as well. Now for part (b), the proof follows from the observation in the previous paragraph since $s_{p}$ is even.
Remark 2.
It is known [19] that
@@
We turn into discussing the case of any arbitrary prime power. In this case, it turns out that $\mathrm{rank}(E(\mathbb{Q}))\leq 1$ provided that $E(\mathbb{Q})[2]$ is non-trivial, see [35, Lemma 2.4] for details. In particular, the Watkins conjecture is true in this case. Furthermore, a complete classification of all such elliptic curves could be found in [35]. The next section will briefly discuss the conjecture for any elliptic curves with a non-trivial $2$-torsion defined over $\mathbb{Q}$.
@@ 3.4. Elliptic curves with non-trivial $2$-torsion
In this section, we assume that $E(\mathbb{Q})[2]$ is either $(\mathbb{Z}/2\mathbb{Z})$ or $(\mathbb{Z}/2\mathbb{Z})^{2}$. One of the important advantages of this assumption is that $2\mid\#E(\mathbb{F}_{p})$ for any prime $p$ co-prime to $2N$, since $E(\mathbb{Q})[2]$ embeds in $E(\mathbb{F}_{p})$. Also, in this case, we have a rank bound in terms of the number of prime factors of $N$. For any squarefree integer $D$, denote $E^{(D)}$ be the twist of $E$ by $D$. Recall that the quadratic twist $E^{(D)}$ is the elliptic curve defined by $y^{2}=x^{3}+aD^{2}x+bd^{3}$. Note that $E^{(D)}$ is isomorphic to $E$ over the field $\mathbb{Q}(\sqrt{D})$. From 1, we have
(10) $\nu_{2}(m_{E^{(D)}})\gg\sum_{\begin{subarray}{c}p\mid D\\ (p,~{}2N)=1\end{subarray}}\nu_{2}\left((p+1)(p+1-a(p))(p+1+a(p))\right)\gg 3(\omega(D)-\omega(N)).$
Then for any $D$ with a sufficiently large number of prime factors, the proof of Watkins’s conjecture for $E^{(D)}$ (see [15]) follows from the following rank bound
(11) $\mathrm{rank}(E(\mathbb{Q}))\leq 2\alpha+\mu-1\leq 2\omega(N)-1,$
@@
Recall that the rank bound at (11) gives the proof of Watkins’s conjecture when $\omega(N)=1$, i.e., when $N$ is a prime power. In fact, in the case when $E$ has a prime power conductor, it is shown in [9] that any quadratic twist $E^{(D)}$ satisfies Watkins’s conjecture. The proof essentially follows from the classification in [35], where a much stronger estimate holds in (10) and (11). More precisely, we have
$\nu_{2}(m_{E^{(D)}})\geq\nu_{2}\left(\frac{m_{E}}{c^{2}_{E}}\right)+\nu_{2}\left(\frac{||f||^{2}_{N^{(D)}}}{||f_{E}||^{2}_{N}}\right)+\frac{1}{6}\nu_{2}\left(\frac{\Delta_{E^{(D)}}}{\Delta_{E}}\right).$
Essentially from Setzer’s classification, the following estimates [9] hold
$\nu_{2}\left(\frac{m_{E}}{c^{2}_{E}}\right)\geq-1,~{}\nu_{2}\left(\frac{||f||^{2}_{N^{(D)}}}{||f_{E}||^{2}_{N}}\right)\gg\omega(D),~{}\nu_{2}\left(\frac{\Delta_{E^{(D)}}}{\Delta_{E}}\right)=O(1),$
@@
In this direction, we also recall the result of Caro and Pasten in [10]. They showed that $E$ satisfies Watkins’s conjecture if $E(\mathbb{Q})[2]=\mathbb{Z}/2\mathbb{Z}$, and all primes of bad reductions, are either non-split or if the number of primes of non-split reduction is odd. In this section, we shall give an extension of this result to the case $E(\mathbb{Q})[2]=(\mathbb{Z}/2\mathbb{Z})^{2}$.
Theorem 6.
@@
Moreover, if $E(\mathbb{Q})[2]=\mathbb{Z}/2\mathbb{Z}$, then Watkins’s conjecture is true for $E$ provided that, either the conditions at [10, Theorem 1.2], or if $\Sha[2]$ is non-trivial.
@@
For the proof of part (a), we follow the same approach as in [10]. If all $E$ has a non-split reduction at every prime, then $\nu_{2}(m_{E})\geq\mu-2$. Suppose that $\mathrm{rank}(E(\mathbb{Q}))=\mu-1$, then it follows from the exact sequence
(12) $0\to E(\mathbb{Q})/2E(\mathbb{Q})\to\mathrm{Sel}^{2}(E)\to\Sha[2]\to 0,$
that $\dim_{\mathbb{F}_{2}}(\mathrm{Sel}^{2}(E))\geq\mu+1$. On the other hand, it immediately follows from the second exact sequence in [10, Lemma 3.1] that
@@
This shows that $\Sha(E)[2]$ is trivial, hence, the parity conjecture holds for $E$, consequently $\mathrm{rank}$ is even, since $-1=-\prod_{p\mid N}w_{p}=w(E/\mathbb{Q})=(-1)^{\mathrm{rank}(E(\mathbb{Q}))}$. This is a contradiction since the rank is $\mu-1$, which is even. Here $w(E/\mathbb{Q})$ denotes the root number of $E/\mathbb{Q}$, and $w_{p}$ denotes the local root number, i.e., the sign of the involution of $W_{p}$ acting at $f_{E}$.
@@
To prove part $(d)$, observe that we have $\nu_{2}(m_{E})\geq\mu-3$ and $\text{rank}(E(\mathbb{Q}))\leq\mu-1$. Now, assume for contradiction that $\text{rank}(E(\mathbb{Q}))>\mu-3$. According to the short exact sequence (12), it follows that $\dim(\mathrm{Sel}^{2}(E))>\mu+1$, which leads to a contradiction. Therefore, the assumption is false, and we can conclude that $\text{rank}(E(\mathbb{Q}))\leq\mu-3$.
Now for the case $E(\mathbb{Q})[2]=\mathbb{Z}/2Z$, it is enough to consider the case when there is at least one split multiplicative reduction. In this case, we have $\nu_{2}(m_{E})\geq\mu-2$. If $\mathrm{rank}(E(\mathbb{Q}))\geq\mu-1$, then we must have $\mathrm{dim}_{\mathbb{F}_{2}}(\Sha[2])=0$. This is because, $\mathrm{dim}_{\mathbb{F}_{2}}(\mathrm{Sel}^{2}(E))\leq\mu$, which follows from [10, Lemma 3.1]. Hence, a contradiction, as $\Sha[2]$ is non-trivial. ∎
Remark 3.
@@
    One can follow a similar treatment for the elliptic curves having additive reductions at some primes. In that case, one needs to put suitable conditions on the parity of $N^{+}-N^{-}$, where $N^{+}$ (resp. $N^{-1}$) denotes the number of $p^{e}$ such that $p^{e}\mid\mid N_{E}$ and $w_{p}=1$ (resp. $w_{p}=-1$). We refer the reader to [27, Theorem 1.1] for a description of the local root numbers at the primes of additive reductions.
@@
    In examining the elliptic curve $E:y^{2}+xy+y=x^{3}+x^{2}-71x-196$, reference to the LMFDB [^2] Refer to https://www.lmfdb.org/EllipticCurve/Q/3315/a/2 for more details. indicates that this curve is semistable and exhibits at least one instance of split multiplicative reduction, alongside possessing a rank of $2$. Further analysis conducted using SageMath reveals that the 2-Selmer rank associated with $E$ is $4$. This set of characteristics ensures that $E$ adheres to the prerequisites outlined in part (b) of Theorem 6.
@@
    For any elliptic curve $E$ mentioned in Theorem 6, one can calculate the $2$-Selmer rank following the methodology outlined in [3, Section 3]. This calculation provides an estimate on the rank of the Shafarevich-Tate group $\Sha[2]$. However, this information is not entirely adequate unless we know the rank of the elliptic curve $E$ over the rational numbers $\mathbb{Q}$.

### v3

@@ 3.3. Elliptic curves with prime conductor
Let $E/\mathbb{Q}$ be any elliptic curve of positive rank, Mestre’s result in [Mestre86] implies that $m_{E}=\langle v_{E},v_{E}\rangle=\sum_{i=1}^{n}w_{i}v_{E}(e_{i})^{2}$. On the other hand, it turns out that $\sum_{i=1}^{n}v_{E}(e_{i})=0$, and all the $w_{i}$ are $1$ when $p=1\pmod{4}$. We immediately have that $m_{E}$ is even in this case. For the case $p=3\pmod{4}$, one needs to use [KK18watkins, Proposition 2.4] and Gross-Kudla’s formula [KK18watkins, Proposition 2.5]. This settles down the conjecture for $\mathrm{rank}(E(\mathbb{Q}))=1$ case. For rank $2$ case, Gross-Kudla’s formula implies $\sum_{i=1}^{n}v_{E}(e_{i})^{3}=0$. The reader may take a look at [KK18watkins] for more details. Consequently we have $\sum_{i=1}^{n}v_{E}(e_{i})^{2}=0\pmod{4}$, if each of the odd terms in $\{v_{E}(e_{i})\}_{1\leq i\leq n}$ appears an even number of times. This is literally true, provided that the following conjecture holds.
@@
When the root number of $E$ is 1-1, Kazalicki and Kohen showed that all such $\nu(e_{i})$ are, in fact, 0. This is true because the root number condition implies $t_{p}(v_{E})=-v_{E}$. For each $1\leq i\leq n$, set $1\leq\bar{i}\leq n$ be the unique index such that $E_{\bar{i}}=E^{p}_{i}$. Note that, $i=\bar{i}$ if and only if, $j(E_{i})\in\mathbb{F}_{p}$. With this, we have [KK17singular] that $t_{p}(e_{i})=e_{\bar{i}}$, which gives a proof of the conjecture for root number 1-1 case. In the root number $1$ case, the conjecture is known to be true if $E$ has a positive discriminant and there exists a prime $\ell$ such that
(3.5) $\left(\frac{-p}{\ell}\right)=-1,~a(\ell)=1\pmod{2}$.
Under these conditions, we have $B_{i,j}(\ell)=0\pmod{2}$, whenever $i,j\in S_{p}$. Moreover, the conditions in (3.5) are satisfied for any $E/\mathbb{Q}$ of conductor $p$, satisfying $E(\mathbb{Q})[2]$ is trivial and $\Delta_{E}>0$.
The main ingredients are $B_{i,j}(\ell)=0\pmod{2}$ and $a(\ell)=1\pmod{2}$. In the case of $\Delta_{E}<0$, the condition $\left(\frac{-p}{\ell}\right)=-1$ implies $\ell$ is inert in $\mathbb{Q}(\sqrt{-p})$ which implies that $2$ divides the order of $\mathrm{Frob}_{\ell}$, which implies that $a(\ell)$ is not odd. This is one of the main hurdles in the case of root number $1$ and negative discriminant. However, Kazalicki and Kohen’s proof also shows that it is enough to have some integer $m$ for which $a(m)=1\pmod{2}$ and all the $B_{i,j}(m)$ are even. From the commutativity of Brandt’s matrices, and [KK17singular, Proposition 20] it follows that $\widetilde{B}(m)=\left(B_{i,j}(n)\right)_{i,j\in S_{p}}\pmod{2}$ is also commutative inside $M_{s_{p}}(\mathbb{Z}/2\mathbb{Z})$. Therefore it may be fruitful to analyze the structure of those $\widetilde{B}(m)$, whose one of the eigenvalues is $1$. Also, note that each $B_{i,j}(m)$ gives a modular form $\theta_{i,j}$, from which we get another modular form $\theta^{\prime}_{i,j}:=\theta_{i,j}-E_{2}$, where $E_{2}=1-24\sum_{n=1}^{\infty}\frac{nq^{n}}{1-q^{n}}$. Now we require the cusp forms $f_{E}$ and $\theta^{\prime}_{i,j}$ to have a common Fourier coefficient of opposite parity. A statement about the independency of the Fourier coefficients, e.g., Generalized Sato-Tate hypothesis [BBG22, Section 4.1], could be helpful in this approach.
We end the discussion of this section with the following remark. As long as the prime conductors are concerned, this is predicted by Watkins in [MW02, Section 4.2] the following.
@@ 3.3.1. A conjecture on the supersingular coefficients
In this section, we discuss the Conjecture 2. In [KK17singular], Kazalicki and Kohen studied modular forms over $\overline{\mathbb{F}_{p}}$ in the sense of Katz’s modular form. If we are a bit more specific, there exists a polynomial $P_{E}(x)\in\mathbb{Z}[x]$ whose roots in $\mathbb{F}_{p}$ are often supersingular zeros. In particular, they showed that if $j\in\mathbb{F}_{p}$ is $j$ invariant of a supersingular elliptic curve $E_{i}$ over $\overline{\mathbb{F}}_{p}$, then $P_{E}(j)=0\pmod{p}$ if and only if, $v_{E}(e_{i})=0\pmod{p}$. In the sense of Katz, the condition $P_{E}(j)=0\pmod{p}$ corresponds to a rule that is 0 acting on $E_{i}$. Finally, a study of Hecke operators on the space of modular forms corresponding to supersingular elliptic curves, again in the sense of Katz. This, of course, does not settle the conjecture; however, this shows that in the case of root number 1-1 case, any supersingular $j$ invariant in $\mathbb{F}_{p}$ is a root of $P_{E}(x)\pmod{p}$. We denote $s_{p}$ to be $S_{p}$, i.e., the number of isomorphism classes of supersingular elliptic curves over $\mathbb{F}_{p}$. Following a remark in Kazalicki-Kohen in [KK17singular], we can deduce the following in support of the conjecture.
Theorem 3.4.
@@
Before proving this, let us first define the required terminology and recall the preliminary results related to the supersingular elliptic curves. For any integer $D=3\pmod{4}$, we say that an elliptic curve over any field $F$ has CM over $\mathcal{O}_{-D}=\mathbb{Z}[\frac{1}{2}(D+\sqrt{-D})]$ if, $\mathcal{O}_{-D}$ contains maximally in the endomorphism ring of the elliptic curve over $\overline{F}$. Then supersingular elliptic curves with $j$-invariant in $\mathbb{F}_{p}$ correspond to those $D$ with those $D$ for which $p$ is ramified or inert in $\mathbb{Q}(\sqrt{-D})$. On the other hand, there are h(D)h(-D) many $\overline{\mathbb{Q}}$ isomorphism classes of elliptic curves over $\overline{\mathbb{Q}}$ with CM by $\mathcal{O}_{-D}$, where h(D)h(-D) is the class number of $\mathbb{Q}(\sqrt{-D})$. Furthermore, we denote $h_{i}(-D)$ to be the number of optimal embeddings of the order of discriminant D-D into $\mathrm{End}(E_{i})$ modulo conjugation by $\mathrm{End}(E_{i})^{*}$ and u(D)u(-D) is the number of units of the order. Then we set $b_{D}=\frac{1}{u(-D)}\sum_{i=1}^{n}h_{i}(-D)v_{E}(e_{i}).$
Proof of Theorem 3.4.
@@
This shows that if there is one even $v_{E}(e_{i})$, then there are at least two even terms since $s_{p}$ is even. Since $\mathrm{rank}(E(\mathbb{Q}))>0$, we have $L(E,1)=0$. In particular, by Gross-Waldspurger formula [Gross85], we have $\langle v_{E},b_{D}\rangle=0$ for any $D>0$ with $\left(\frac{-D}{p}\right)=-1$.
For proof of part (a), note that since $p$ and $D$ satisfy the imposed conditions, $p$ is inert in $\mathbb{Q}(\sqrt{-D})$. In particular, $b_{D}$ is supported only on the indices in $S_{p}$. Moreover, since $\mathbb{Q}(\sqrt{-D})$ has class number $1$, we have $h_{i}(D)=1$ for exactly one $i\in S_{p}$, and that $E_{i}$ has CM by $\mathcal{O}_{-D}$. Then the condition $\langle v_{E},b_{D}\rangle=0$ implies $v_{E}(e_{i})=0$. In particular, if $s_{p}=2$, then $v_{E}(e_{i^{\prime}})=0$ for the other $i^{\prime}\in S_{p}$, as well. Now for part (b), the proof follows from the observation in the previous paragraph since $s_{p}$ is even.
Remark 3.5.
It is known [Gross85] that
@@
We turn into discussing the case of any arbitrary prime power. In this case, it turns out that $\mathrm{rank}(E(\mathbb{Q}))\leq 1$ provided that $E(\mathbb{Q})[2]$ is non-trivial, see [Setzer75, Lemma 2.4] for details. In particular, the Watkins conjecture is true in this case. Furthermore, a complete classification of all such elliptic curves could be found in [Setzer75]. The next section will briefly discuss the conjecture for any elliptic curves with a non-trivial $2$-torsion defined over $\mathbb{Q}$.
@@ 3.4. Elliptic curves with non-trivial $2$-torsion
In this section, we assume that $E(\mathbb{Q})[2]$ is either $(\mathbb{Z}/2\mathbb{Z})$ or $(\mathbb{Z}/2\mathbb{Z})^{2}$. One of the important advantages of this assumption is that $2\mid\#E(\mathbb{F}_{p})$ for any prime $p$ co-prime to $2N$, since $E(\mathbb{Q})[2]$ embeds in $E(\mathbb{F}_{p})$. Also, in this case, we have a rank bound in terms of the number of prime factors of $N$. For any squarefree integer $D$, denote $E^{(D)}$ be the twist of $E$ by $D$. Recall that the quadratic twist $E^{(D)}$ is the elliptic curve defined by $y^{2}=x^{3}+aD^{2}x+bD^{3}$. Note that $E^{(D)}$ is isomorphic to $E$ over the field $\mathbb{Q}(\sqrt{D})$. From (2.1), we have
(3.6) $\nu_{2}(m_{E^{(D)}})\gg\sum_{\begin{subarray}{c}p\mid D\\ (p,~2N)=1\end{subarray}}\nu_{2}\left((p+1)(p+1-a(p))(p+1+a(p))\right)\gg 3(\omega(D)-\omega(N)).$
Then for any $D$ with a sufficiently large number of prime factors, the proof of Watkins’s conjecture for $E^{(D)}$ (see [EP21]) follows from the following rank bound
(3.7) $\mathrm{rank}(E(\mathbb{Q}))\leq 2\alpha+\mu-1\leq 2\omega(N)-1,$
@@
Recall that the rank bound at (3.7) gives the proof of Watkins’s conjecture when $\omega(N)=1$, i.e., when $N$ is a prime power. In fact, in the case when $E$ has a prime power conductor, it is shown in [caro22watkinss] that any quadratic twist $E^{(D)}$ satisfies Watkins’s conjecture. The proof essentially follows from the classification in [Setzer75], where a much stronger estimate holds in (3.6) and (3.7). More precisely, we have
[…]
Essentially from Setzer’s classification, the following estimates [caro22watkinss] hold
[…]
@@
In this direction, we also recall the result of Caro and Pasten in [CP22]. They showed that $E$ satisfies Watkins’s conjecture if $E(\mathbb{Q})[2]=\mathbb{Z}/2\mathbb{Z}$, and all primes of bad reductions, are either non-split or if the number of primes of non-split reduction is odd. In this section, we shall give an extension of this result to the case $E(\mathbb{Q})[2]=(\mathbb{Z}/2\mathbb{Z})^{2}$.
Theorem 3.6.
@@
Moreover, if $E(\mathbb{Q})[2]=\mathbb{Z}/2\mathbb{Z}$, then Watkins’s conjecture is true for $E$ provided that, either the conditions at [CP22, Theorem 1.2], or if $\Sha[2]$ is non-trivial.
@@
For the proof of part (a), we follow the same approach as in [CP22]. If all $E$ has a non-split reduction at every prime, then $\nu_{2}(m_{E})\geq\mu-2$. Suppose that $\mathrm{rank}(E(\mathbb{Q}))=\mu-1$, then it follows from the exact sequence
(3.8) $0\to E(\mathbb{Q})/2E(\mathbb{Q})\to\mathrm{Sel}^{2}(E)\to\Sha[2]\to 0,$
that $\dim_{\mathbb{F}_{2}}(\mathrm{Sel}^{2}(E))\geq\mu+1$. On the other hand, it immediately follows from the second exact sequence in [CP22, Lemma 3.1] that
@@
This shows that $\Sha(E)[2]$ is trivial, hence, the parity conjecture holds for $E$, consequently $\mathrm{rank}$ is even, since $1=-\prod_{p\mid N}w_{p}=w(E/\mathbb{Q})=(-1)^{\mathrm{rank}(E(\mathbb{Q}))}$. This is a contradiction since the rank is $\mu-1$, which is even. Here $w(E/\mathbb{Q})$ denotes the root number of $E/\mathbb{Q}$, and $w_{p}$ denotes the local root number, i.e., the sign of the involution of $W_{p}$ acting at $f_{E}$.
@@
To prove part $(d)$, observe that we have $\nu_{2}(m_{E})\geq\mu-3$ and $\text{rank}(E(\mathbb{Q}))\leq\mu-1$. Now, assume for contradiction that $\text{rank}(E(\mathbb{Q}))>\mu-3$. According to the short exact sequence (3.8), it follows that $\dim(\mathrm{Sel}^{2}(E))>\mu+1$, which leads to a contradiction. Therefore, the assumption is false, and we can conclude that $\text{rank}(E(\mathbb{Q}))\leq\mu-3$.
Now for the case $E(\mathbb{Q})[2]=\mathbb{Z}/2Z$, it is enough to consider the case when there is at least one split multiplicative reduction. In this case, we have $\nu_{2}(m_{E})\geq\mu-2$. If $\mathrm{rank}(E(\mathbb{Q}))\geq\mu-1$, then we must have $\mathrm{dim}_{\mathbb{F}_{2}}(\Sha[2])=0$. This is because, $\mathrm{dim}_{\mathbb{F}_{2}}(\mathrm{Sel}^{2}(E))\leq\mu$, which follows from [CP22, Lemma 3.1]. Hence, a contradiction, as $\Sha[2]$ is non-trivial. ∎
Remark 3.7.
@@
    One can follow a similar treatment for the elliptic curves having additive reductions at some primes. In that case, one needs to put suitable conditions on the parity of $N^{+}-N^{-}$, where $N^{+}$ (resp. $N^{-1}$) denotes the number of $p^{e}$ such that $p^{e}\mid\mid N_{E}$ and $w_{p}=1$ (resp. $w_{p}=-1$). We refer the reader to [rootnumber2000, Theorem 1.1] for a description of the local root numbers at the primes of additive reductions.
@@
    In examining the elliptic curve $E:y^{2}+xy+y=x^{3}+x^{2}-71x-196$, reference to the LMFDB [^2] Refer to https://www.lmfdb.org/EllipticCurve/Q/3315/a/2 for more details. indicates that this curve is semistable and exhibits at least one instance of split multiplicative reduction, alongside possessing a rank of $2$. Further analysis conducted using SageMath reveals that the 2-Selmer rank associated with $E$ is $4$. This set of characteristics ensures that $E$ adheres to the prerequisites outlined in part (b) of Theorem 3.6.
@@
    For any elliptic curve $E$ mentioned in Theorem 3.6, one can calculate the $2$-Selmer rank following the methodology outlined in [RPC, Section 3]. This calculation provides an estimate on the rank of the Shafarevich-Tate group $\Sha[2]$. However, this information is not entirely adequate unless we know the rank of the elliptic curve $E$ over the rational numbers $\mathbb{Q}$.